OpenStudy (anonymous):

If x2 + y2 = 100, and xy = 18, find the value of x - y.

OpenStudy (anonymous):

8

OpenStudy (akshay_budhkar):

see x2+y2-2xy=(x-y)2.. you can substitute the values to get your answer

OpenStudy (anonymous):

wrong

OpenStudy (angela210793):

just express x=18/y and replace it in the other eq

OpenStudy (akshay_budhkar):

the answer is 8

OpenStudy (akshay_budhkar):

8 is right

OpenStudy (anonymous):

there are two answers.

OpenStudy (akshay_budhkar):

+8,-8 then..

OpenStudy (anonymous):

yes right

OpenStudy (akshay_budhkar):

so you get it how to solve?

OpenStudy (anonymous):

im on my next problem now

OpenStudy (akshay_budhkar):

okay. you get this right no?

myininaya (myininaya):

(x-y)^2=x^2-2xy+y^2=x^2+y^2-2xy=(100)-2(18)=100-36= 64 take square root of both sides x-y= plus or minus 8